As of January 1, Year 2, Concepts Inc. had a balance of $4,500 in Cash and $2,500 in Common Stock, and $2,000 in Retained Earnings. These were the only accounts with balances at the beginning of Year 2.

Further analysis of the company’s cash account indicated that during Year 2 the company had (1) net cash inflow from operating activities of $5,100, (2) net cash outflow for investing activities of $13,000, and (3) net cash inflow from financing activities of $7,600. All revenue and expense events were cash events.

Concepts Inc. ended Year 2 with the following accounts and balances:

  • Cash 4,200
  • Land 13,000
  • Notes Payable 3,000
  • Common Stock 8,000
  • Retained Earnings 2,000
  • Revenue 9,900
  • Expenses 4,800
  • Dividends 900

1. The Land account had a beginning balance of zero and and ending balance of $13,000. If this change was due to one transaction, what could have accounted for this change?

2. The Notes Payable account had a beginning balance of zero and and ending balance of $3,000. If this change was due to one transaction, what could have accounted for this change?

3. The Common Stock account had a beginning balance of $2,500 and and ending balance of $8,000. If this change was due to one transaction, what could have accounted for this change?

4. How much net income did Concepts Inc. realize for Year 2?

5. At the end of Year 2, both sides of the accounting equation equal:
Answer #1
Journal Entry:
Accounts Title Debit $ Credit$
1) The transaction would be : Land 13000
"Purchase of land partially Cash 10000 & issue Cash 10000
of Notes 3000, amounting $13000" Notes Payable 3000
2) The transaction would be : Land 3000
"Issue of Notes for land, amounting $3000" Notes Payable 3000
3) The transaction would be : Cash 5500
"Issue of Common Shares for cash, amounting Common Stock 5500
$5500"
4) Net Income realized for Year 2: Revenue - expenses = 9900-4800 = 5100
5) Both sides of accounting equation:
Assets side = Cash 4200 + Land 13000 = 17200
Liabilities side = Notes payable 3000+ Common Stock 8000 + Retained Earnings (2000 + 5100- 900) = 17200